Spring 2019 FIN 341 Case Study on Capital Budgeting Hailey just started her first job at Whatcom Co. as a junior budget analyst. She is working for the Venture Capital Division and has been given for capital budgeting projects to evaluate. She must give her analysis and recommendation to the capital budgeting committee. Hailey has a B.S. in accounting from WWU (2007) and passed the CPA exam (2008). She has been in public accounting for 2 years. During that time, she earned an MBA from Seattle U. She would like to be the CFO of a company someday--maybe Whatcom Co. -- and this is an opportunity to get onto that career track and to prove her ability. As Hailey looks over the financial data collected, she is trying to make sense of it all. She already has the most difficult part of the analysis complete -- the estimation of cash flows. Through some internet research and application of finance theory, she has also determined the firm’s beta. Here is the information that Hailey has accumulated so far: The Capital Budgeting Projects She must choose one of the four capital budgeting projects listed below: Table 1 Table 1 shows the expected after-tax operating cash flows for each project. All projects are expected to have a 4-year life. The projects differ in size (the cost of the initial investment), and their cash flow patterns are different. They also differ in risk as indicated in the above table. The capital budget is $20 million, and the projects are mutually exclusive. Capital Structures Whatcom Co. has the following capital structure, which is considered to be optimal: Common Equity 50% Preferred Equity 10% Debt 40% 100% Cost of Capital Hailey knows that in order to evaluate the projects she will have to determine the cost of capital for each of them. She has been given the following data, which he believes will be relevant to her task.

(1) The firm’s tax rate is 40%.

(2) Whatcom Co. has issued a 12% semi-annual coupon bond with 5 years term to maturity. The current trading price is $1080.

(3) The firm has issued some preferred stock which pays an annual 11% dividend of $100 par value, and the current market price is $104.

(4) The firm’s stock is currently selling for $94 per share. Its last dividend (D0) was $3, and dividends are expected to grow at a constant rate of 9%. The current risk-free return offered by Treasury security is 2.5%, and the market portfolio’s return is 10%. Whatcom Co. has a beta of 1.3

(5) The firm adjusts its project WACC for risk by adding 2.5% to the overall WACC for high-risk projects and subtracting 2.5% for low-risk projects. Hailey knows that Whatcom Co. executives have favored IRR in the past for making their capital budgeting decisions.

Her professor at Seattle U. said NPV was better than IRR. She is the new kid on the block and must be prepared to defend her recommendations. First, however, Hailey must finish the analysis and write her report. To help begin, she has formulated the following questions:

1. What is the firm’s cost of debt?

2. What is the cost of preferred stock for Whatcom Co.?

3. Cost of common equity (1) What is the estimated cost of common equity using the CAPM approach? (2) What is the estimated cost of common equity using the DCF approach? (3) What is the final estimate for cost of equity?

4. What is Whatcom Co.’s overall WACC?

5. Do you think the firm should use the single overall WACC as the hurdle rate for each of its projects? Explain.

6. What is the WACC for each project?

7. Calculate all relevant capital budgeting measures for each project and place your numerical solutions in Table 2. Table 2 A B C D NPV IRR

8. Comment on the commonly used capital budgeting measures. What is the underlying cause of ranking conflicts? Which criterion is the best one, and why?

9. Which of the projects are unacceptable and why?

10. Rank the projects that are acceptable, according to Hailey’s criterion of choice.
Answer #1

You have asked 10 questions in a single post. I will address first four of them. Please post the balance questions separately.

1. What is the firm’s cost of debt?

(Pre tax) Cost of debt, Kd = YTM of the bond = 2 x RATE (Period, PMT, PV, FV) = 2 x RATE (2 x 5, 12%/2 x 1000, -1080, 1000) = 2 x RATE (10, 60, -1080, 1000) = 9.93%

2. What is the cost of preferred stock for Whatcom Co.?

Ks = Annual dividend / Current price = 11% x 100 / 104 = 10.58%

3. Cost of common equity

(1) What is the estimated cost of common equity using the CAPM approach?

Ke = Rf + Beta x (Rm - Rf) = 2.5% + 1.3 x (10% - 2.5%) = 12.25%

(2) What is the estimated cost of common equity using the DCF approach?

Ke = D0 x (1 + g) / P + g = 3 x (1 + 9%) / 94 + 9% = 12.48%

(3) What is the final estimate for cost of equity?

We can take average of the two, Ke = (12.25% + 12.48%) / 2 = 12.36%

4. What is Whatcom Co.’s overall WACC?

WACC = Wd x Kd x (1 - T) + Ws x Ks + We x Ke = 40% x 9.93% x (1 - 40%) + 10% x 10.58% + 50% x 12.36% = 9.62%

Please post the balance questions separately.